OpenStudy (anonymous):

Potassium hyroxide and phosphoric acid react to form potassium phosphate and water according to the equation: 3KOH + H3PO4 --> K3PO4 + 3H2O Determine the starting mass of each reactant if 63.5 g of K3PO4 is produced and 84.8 g of H3PO4 remains unreacted. What is the starting mass of KOH? What is the starting mass of H3PO4?

OpenStudy (anonymous):

Alright so we are going to ignore the mass that did not react for now and use the nass that did react to determine how much KOH it took to cause the reaction.

OpenStudy (anonymous):

Wait i read that wrong let me make some quick rearrangements lol

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

Okay so we are going to use the amount of KOH produced to determine how much H3PO4 reacted.

OpenStudy (anonymous):

*K3PO4

OpenStudy (anonymous):

So 63.5g

OpenStudy (anonymous):

yeah im doing my work right now but what we want to do is start with the 63.5 g of K3PO4 and convert it to KOH in g

OpenStudy (anonymous):

\[\frac{ 63.5 g K3PO4 }{ }*\frac{ 1 mole K3PO4 }{ 212 g K3PO4 }*\frac{ 3 mole KOH }{ 1 mole K3PO4 }*\frac{ 56 kg KOH }{ 1 mole KOH }*\frac{ 1000 g }{ 1 kg }\]

OpenStudy (anonymous):

that would give you the starting g of KOH. and you do the same thing for H3PO4 but you will add 84.58 to whatever you get for H3PO4 because that amount did not react. And the conversion give you the amount that did react.

OpenStudy (anonymous):

I got 50320.754 for KOH

OpenStudy (anonymous):

What would I use with the H3PO4? Would it be the H2O?

OpenStudy (anonymous):

KOH is 50.32 i messed up on the conversion lol see i used Kg thats supposed to be gram so ignore that last step where you multiplied by 1000. And no H2O is just bonus information you dont really need it for H3PO4 use K3PO4 also its just that the ratios will be different when its mole over mole.

OpenStudy (anonymous):

Do you think you could set up the one for H3PO4 or do you want me to do it for you

OpenStudy (anonymous):

I'll try it

OpenStudy (anonymous):

Alright ill be here

OpenStudy (anonymous):

84.8 g x (1 mol H3PO4/ 97.994) x ( 1 mol K3PO4/ 1 mol H3PO4) x ( 56.10g K3PO4 / mol H3PO4)

OpenStudy (anonymous):

No , You want to find out how much H3PO4 it took to create 56.1 g of K3PO4. To do this you want to find out how much H3PO4 is made from 56.1 g of K3PO4. STart with 56.1 g of K3PO4 and convert it to H3PO4 in grams.

OpenStudy (anonymous):

and its 63.5 not 56.1

OpenStudy (anonymous):

Sorry

OpenStudy (anonymous):

no its fine just remember to basically do the problem backwards if youre looking for H3PO4 start with the K3PO4

OpenStudy (anonymous):

\[\frac{ 63.5 g K3PO4 }{ }*\frac{ 1 mole K3PO4 }{ 212 g K3PO4 }*\frac{ 1 mole H3PO4 }{ 1 mole K3PO4 }*\frac{ 98 g H3PO4 }{ 1 mole H3PO4 }\]

OpenStudy (anonymous):

so would it basically be like the K3PO4 except replace the KOH with H3PO4?

OpenStudy (anonymous):

yes that is right. Youre just switching what youre converting to so the only thing that changes is the molar mass and possibly the ratio of moles

OpenStudy (anonymous):

that will give you the amount of H3PO4 that reacted.

OpenStudy (anonymous):

29.35377

OpenStudy (anonymous):

yes so what it the starting mass of H3PO4 and KOH?

OpenStudy (anonymous):

You need to add 84.8 to it right?

OpenStudy (anonymous):

Yes! I wanted to make sure you understood that!

OpenStudy (anonymous):

So I got 114.153

OpenStudy (anonymous):

Yes that is the answer for K3PO4, what about for KOH?

OpenStudy (anonymous):

50.32

OpenStudy (anonymous):

Yes Sir!

OpenStudy (anonymous):

Nice

OpenStudy (anonymous):

Thanks for the help. I wish I could give you extra medals. You've really helped me today.

OpenStudy (anonymous):

yes sir anything else?

OpenStudy (anonymous):

Not right now. I'm going through right now making sure I didn't skip anything.

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

oh and you're welcome! lol and it is cool i use this as a break from my homework it is fun!

OpenStudy (anonymous):

What's your homework?

OpenStudy (anonymous):

And you would make a really good chemistry tutor if you're not already. I know my school pays people to tutor.

OpenStudy (anonymous):

Its Calculas but more specifically optimization right now lol. and thank you lol and i would tutor but im a cashier at dollar tree :P

OpenStudy (anonymous):

Yikes. I avoided that stuff like the plague. Then again I'm not the best at math.

OpenStudy (anonymous):

Yeah lol i do not like it either! just one more semester of math after this and im done with it for life lol

OpenStudy (anonymous):

I thought that too but then learned how much math is in chemistry then I got physics too. Bleh. I got my actual math classes out of the way though. Do you have a major?

OpenStudy (anonymous):

omg i forgot about physics. but I have never taken a physics class in my life so that might make it harder too lol. And yeah im a biochemistry major what about you?

OpenStudy (anonymous):

I haven't taken a physic course a day in my life either. I'm a pre-vet major.

OpenStudy (anonymous):

to be honest i don't even know what Phsyics is lol and im a pre-med